OpenStudy (anonymous):

Help! "Show that the Tension at the bottom of any pendulum will be equal to 3 times the weight of the pendulum. Show all steps, equations and free body diagrams." This is for my lab due tom and i dont understand how to do this!

OpenStudy (anonymous):

@kurodin which type of pendulum is in your Lab?simple or compound pendulum...

OpenStudy (anonymous):

ok for any pendulum..

OpenStudy (anonymous):

\[(m \times v ^{2})/l+m \times g=T\]....(i) from the conservation of energy:: \[(1\div 2)m \times v ^{2}=m \times g \times l\]...(ii) solve (i) and (ii) T=3\[3\times m \times g\] hence it will be 3 times to the weight...

OpenStudy (anonymous):